# HOSIG: Full-Body Human-Object-Scene Interaction Generation with Hierarchical Scene Perception

### v1

@@ Abstract.
Generating high-fidelity full-body human interactions with dynamic objects and static scenes remains a critical challenge in computer graphics and animation. Existing methods for human-object interaction often neglect scene context, leading to implausible penetrations, while human-scene interaction approaches struggle to coordinate fine-grained manipulations with long-range navigation. To address these limitations, we propose HOSIG, a novel framework for synthesizing full-body interactions through hierarchical scene perception. Our method decouples the task into three key components: 1) a scene-aware grasp pose generator that ensures collision-free whole-body postures with precise hand-object contact by integrating local geometry constraints, 2) a heuristic navigation algorithm that autonomously plans obstacle-avoiding paths in complex indoor environments via compressed 2D floor maps and dual-component spatial reasoning, and 3) a scene-guided motion diffusion model that generates trajectory-controlled, full-body motions with finger-level accuracy by incorporating spatial anchors and dual-space classifier-free guidance. Extensive experiments on the TRUMANS dataset demonstrate superior performance over state-of-the-art methods. Notably, our framework supports unlimited motion length through autoregressive generation and requires minimal manual intervention. This work bridges the critical gap between scene-aware navigation and dexterous object manipulation, advancing the frontier of embodied interaction synthesis. Codes will be available after publication. Project page: https://yw0208.github.io/hosig/
Character-scene interaction, motion synthesis, scene representation
submissionid: 382ccs: Computing methodologies Computer graphics Animation
@@
Figure 1. Human-Object-Scene Interaction Generation. Our proposed HOSIG can generate high-fidelity full-body human motions. HOSIG can not only generate interactions with static scenes, but also generate object manipulation motions with fine hand-object contact. Moreover, relying on iterative generation and autonomous navigation, HOISG can generate long-term motions in complex indoor scenes.
1. Introduction
@@
Embodied intelligence represents a pivotal frontier in AI research, aiming to develop agents capable of navigating and manipulating 3D environments. While existing works have extensively explored human-object interaction (HOI) (Taheri et al., 2024; Zhang et al., 2024a; Petrov et al., 2024; Wu et al., 2022; Zheng et al., 2023; Li et al., 2023) and human-scene interaction (HSI) (Karunratanakul et al., 2023; Xiao et al., 2023; Zhao et al., 2023; Zhang and Tang, 2022a; Wang et al., 2024a), few address the integrated human-object-scene interaction (HOSI) challenge (Wu et al., 2024; Jiang et al., 2024b, a; Lu et al., 2024). As shown in Figure 1, our goal is to enable characters to move in complex scenes, complete precise object operations, and seamlessly connect them into a long-term motion. The inherent complexity of synthesizing these multimodal interactions presents critical unsolved challenges in spatial reasoning and motion coordination.
@@
While recent advances in HOI (Yang et al., 2024; Song et al., 2024; Peng et al., 2023; Taheri et al., 2022; Ghosh et al., 2023; Zhang et al., 2024b), they predominantly neglect 3D scene constraints. This oversight leads to implausible human-scene interpenetration, highlighting the necessity for scene-aware reasoning. Besides ignoring the scene in HOI, current HSI approaches exhibit two limitations: 1) Global scene encoding methods (Wang et al., 2024a, 2022a; Huang et al., 2023) lack granularity for precise motion synthesis, and 2) Local context perception strategies (Cen et al., 2024; Mao et al., 2022; Ghosh et al., 2021) struggle with pathfinding in complex environments, resulting in persistent interpenetration during long-range motion generation. These dual challenges underscore the need for unified scene-object-agent coordination.
@@
Our technical implementation features three key innovations. First, the grasp pose generator augments the cVAE framework (Taheri et al., 2022) with local scene geometry constraints, producing physically-plausible hand-object orientations that prevent scene interpenetration. Second, a novel 2D scene abstraction layer enables efficient navigation through 3D environments via heuristic pathfinding on compressed obstacle-aware maps, dynamically generating obstacle-aware motion graphs. Third, building upon ControlNet’s conditioning paradigm (Zhang et al., 2023), we develop a multi-condition diffusion framework that simultaneously integrates: (a) spatial anchors from navigation paths, (b) fine-grained hand control through grasp poses, and (c) path-aligned scene priors for continuous interaction optimization. This unified architecture achieves finger-level motion precision without auxiliary hand modules, surpassing previous trajectory-control methods through iterative spatial constraint.
@@
2. Related Work
2.1. Human-Object Interaction Generation
Research on human-object interaction (HOI) motion generation has progressed through two main paradigms. Early approaches primarily utilized reinforcement learning (RL), with initial works like (Peng et al., 2019, 2021) achieving basic interactions such as box touching. Subsequent RL methods developed more complex skills including basketball dribbling (Liu and Hodgins, 2018), tennis playing (YUAN and Makoviychuk, 2023), multi-object manipulation (Wang et al., 2023), and box carrying (Hassan et al., 2023).
The second paradigm employs generative models for direct motion synthesis. GOAL (Taheri et al., 2022) pioneered this direction using cVAEs for grasping motions. IMoS (Ghosh et al., 2023) later enabled action-label conditioned generation (e.g., photography, drinking). Recent works focus on language-conditioned generation: OOD-HOI (Zhang et al., 2024a) synthesizes human-object motions separately with contact optimization, while HOI-Diff (Peng et al., 2023) employs affordance-guided diffusion. Other key innovations include TriDi’s unified modeling of bodies/objects/contacts (Petrov et al., 2024) and CHOIS’s trajectory-constrained synthesis (Li et al., 2024a). InterDiff (Xu et al., 2023) further enables sequential generation from initial states.
2.2. Human-Scene Interaction Generation
@@
Figure 2. Overview of Our Pipeline. HOSIG can iteratively generate long-term motions based on spatial information, text, and the previous motion clip. There are three parts worth noting in the pipeline: (1) SGAP (Section 3.2) generates fine grasping postures to ensure the quality of character interaction. (2) Heuristic Navigation (Section 3.3) generates sparse human root joint trajectories to constrain the subsequent generated motions to be within the traversable area. (3) SCoMoGen (Section 3.4) uses a dual-branch design to achieve spatial control and adds additional joint & scene guidance during inference to achieve high-precision control.
For locomotion generation, existing methods adopt either data-driven (Wang et al., 2024b; Zhang and Tang, 2022b; Rempe et al., 2023) or algorithmic approaches (Wang et al., 2022b). While data-driven methods struggle in novel complex environments, our work proposes an algorithmic solution for robust navigation. Semantic interaction synthesis initially produced static poses in scenes (Zhang et al., 2020; Li and Dai, 2024; Zhao et al., 2022; Xuan et al., 2023), later evolving into language-guided systems (Wang et al., 2022a, 2024a; Huang et al., 2023) though requiring post-hoc optimization for scene compliance. Recent advances include concurrent navigation-interaction frameworks (Yi et al., 2024) and video generation-based methods (Li et al., 2024b) with strong generalization but limited contact realism and motion range.
Closest to our approach, (Jiang et al., 2024b, a) enable triadic human-object-scene interactions. However, these require frame-level action specifications and precise finger motion guidance, whereas our method achieves comparable results using only initial/final object states.
3. Method
3.1. Problem Formulation
@@
3.2. Scene-Aware Grasp Pose Generation
@@
(1) $X=\left[\varTheta,\beta,\mathbf{V},\mathbf{D}^{h\rightarrow o},\hat{\mathbf{h}},\mathbf{t}^{o},\mathbf{B}^{o},\mathbf{B}^{s}\right]$
@@
Figure 3. Visulization of Scene-Aware in SGAP. The green box is centered on the purple bottle. In SGAP, only the sparse scene point cloud inside the box is used, as shown in the right figure.
The critical component $\mathbf{B}^{s}$ constitutes the primary scene perception mechanism in SGAP. As illustrated in Figure 3, given an object translation $\mathbf{t}^{o}=(x,y,z)$, we construct a scene context volume bounded by $[x-0.8,x+0.8]\times[y-0.8,y+0.8]\times[0.2,1.8]$ meters, forming a $1.6^{3}\ m^{3}$ cubic region. To capture scene geometry, we employ a volumetric sampling strategy that generates a dense point cloud $\mathcal{P}_{s}$ containing both surface vertices and interior points from the scene mesh. This sampling follows a voxel grid resolution of 8 cm³, ensuring complete spatial coverage. The BPS transformation processes $\mathcal{P}_{s}$ through basis projections to produce the scene encoding $\mathbf{B}^{s}\in\mathbb{R}^{1024}$. These scene features condition the pose generation process, enabling synthesis of scene-adapted full-body poses.
@@
(2) $\mathcal{L}_{sd}=\frac{1}{N}\sum_{j=1}^{22}\sum_{k=1}^{N}\|\mathbf{J}_{j}-\mathbf{S}_{k}\|_{2}^{2}$
@@
where $\mathbf{J}_{j}\in\mathbb{R}^{3}$ denotes the $j$-th body joint position from the SMPL-X kinematic tree, $\mathbf{S}_{k}\in\mathbb{R}^{3}$ represents the $k$-th point in the localized scene point cloud $\mathcal{P}_{s}$ with variable point numbers $N$. This formulation imposes progressively stronger penalties as joints approach scene surfaces, effectively creating a repulsive potential field. During training, $\mathcal{L}_{sd}$ backpropagates collision-avoidance constraints through the cVAE’s latent space, encouraging the generator to produce poses maintaining clearance from scene geometry while preserving natural motion kinematics.
@@
During inference, SGAP’s output parameters undergo selective distillation to extract three critical control signals: root joint translation $\mathbf{a}_{r}\in\mathbb{R}^{3}$, wrist joint position $\mathbf{a}_{w}\in\mathbb{R}^{3}$, and hand joint rotations $\boldsymbol{\theta}_{h}\in\mathbb{R}^{15\times 6}$. These elements constitute the spatial anchor tuple $\mathcal{A}=(\mathbf{a}_{r},\mathbf{a}_{w},\boldsymbol{\theta}_{h})$, which serves as key constraints for subsequent motion synthesis.
Input: 3D scene point cloud $\mathcal{S}_{p}$, trajectory keypoints $\{P_{i}\}_{i=0}^{n}$.
Output: Optimized sparse path $\{path_{i}\}_{i=0}^{n-1}$.
1
2Initialize: Smooth 2D obstacle-aware map $\mathcal{S}_{map}\leftarrow\mathcal{S}_{p}$, normalized start/goal node pairs $\{node^{s}_{i},node^{e}_{i}\}_{i=0}^{n-1}\leftarrow\{P_{i}\}_{i=0}^{n}$.
3for each $(node^{s}_{i},node^{e}_{i})$ in $\{node^{s}_{i},node^{e}_{i}\}_{i=0}^{n-1}$ do
4
5     Initialize priority queue $\mathcal{Q}$, g-score table $\mathcal{G}$, came-from table $\mathcal{C}$;
6     Compute f-score $\mathcal{F}(node^{s}_{i})\leftarrow heuristic(node^{s}_{i},\ node^{e}_{i})$;
7     Add $\{node^{s}_{i}:\mathcal{G}(node^{s}_{i})+\mathcal{F}(node^{s}_{i})\}$ to $\mathcal{Q}$;
8     while $\mathcal{Q}$ is not empty do
9           Extract current node $node^{q}$ with minimal value from $\mathcal{Q}$;
10          if $current=goal$ then
11                Backtrack and record dense path $path$;
12               break;
13                end if
14
15               for 4-directional $neighbors$ of $node^{q}$ do
16                     if new path is better then
17                          Update $\mathcal{G}$, $\mathcal{F}$, $\mathcal{C}$, and push to $\mathcal{Q}$;
18                          end if
19
20                          end for
21
22                          end while
23
24                         Adjust path height and coordinates to original space;
25                         Save $path_{i}$;
26                          end for
ALGORITHM 1 Heuristic Navigation Algorithm
3.3. Heuristic Navigation on 2D Obstacle-Aware Map
Unlike learning-based methods that require extensive training data (Yi et al., 2024), Algorithm 1 operates on 3D scene pointcloud $\mathcal{S}_{p}$ while maintaining computational efficiency. Next, we will introduce several key points in the algorithm.
The 3D-to-2D projection converts $S_{p}$ into a navigable obstacle-aware map $\mathcal{S}_{map}=\{v_{ij}\}$ via vertical voxel quantization. Let $G=\{g_{ij}\}$ denote the 2D grid where each cell $g_{ij}$ corresponds to a $0.1\text{m}\times 0.1\text{m}$ planar region. The elevation value $v_{ij}$ of $g_{ij}$ is computed as:
[…]
where $\mathbb{I}$ is the indicator function. In short, Equation 3 is to calculate the number of scene mesh vertices within the range of $g_{ij}$. As for $20.2<z<$ 2, this is to avoid counting useless information such as the floor and ceiling. This can effectively prevent the character from walking against obstacles. This vertical point density encoding effectively captures obstacles while preserving walkable surfaces, as demonstrated in Figure 1 and 2, red means obstacles ($v_{ij}>0$) and green means walkable ($v_{ij}=0$).
Our dual-component heuristic $\mathcal{F}$ combines geometric and density awareness. For any two nodes $a=(x_{a},y_{a})$ and $b=(x_{b},y_{b})$, we define:
(4) $\mathcal{F}(a,b)=\underbrace{\|a-b\|_{2}}_{\text{Euclidean term}}+\lambda\sum_{g\in\mathcal{L}(a,b)}v_{g}$
where $\mathcal{L}(a,b)$ denotes the Bresenham line algorithm’s output between $a$ and $b$, and $\lambda$ controls obstacle avoidance strength. This formulation guides the search away from obstacle areas while maintaining acceptable arrival efficiency.
As shown in Algorithm 1, the “new path is better” condition evaluates both accumulated cost and future estimates. Let $\mathcal{G}(node^{q})$ denote the current path cost from $node^{s}$ to node $node^{q}$. A candidate neighbor $node^{q^{\prime}}$ replaces existing entries in $\mathcal{G}$ if:
(5) $\mathcal{G}(q)+v_{node^{q^{\prime}}}<\mathcal{G}(node^{q^{\prime}})$
which means a better path (from $node^{s}$ to $node^{q}$ to $node^{q^{\prime}}$) is found. Meanwhile, make $\mathcal{G}(node^{q^{\prime}})=\mathcal{G}(q)+v_{node^{q^{\prime}}}$, $\mathcal{F}$ records the heuristic value from $node^{q^{\prime}}$ to the end node $node^{e}$, $\mathcal{C}$ updates $\mathcal{C}(node^{q^{\prime}})=node^{q}$ and $\mathcal{Q}$ updates $\{node^{q}:\mathcal{G}(node^{q})+\mathcal{F}(node^{q})\}$. In the next iteration, current node $node^{q}$ is the node with the minimal value, because $\mathcal{Q}$ is a priority queue. Just keep going forward until reaching the end point $node^{e}$. When the $node^{q}$ is the end point, trace back the path from the came-from table $\mathcal{C}$.
In actual use, we can predefine or generate (i.e., $a_{r}$ generated by SGAP) any number of ordered nodes, and our algorithm can obtain a feasible path to travel through the scene. For convenience, the path is written as $a_{r}$ just like $a_{r}$ generated by SGAP.
3.4. Scene-Guided Controllable Motion Generation
@@
To enable the generation of high-fidelity interactions, two key challenges must be addressed. First, effective control over human body movement along the specified trajectory is required, with particular emphasis on enabling precise hand movements for object grasping. Second, in narrow and complex scenes, relying solely on the trajectory is insufficient to guarantee scene avoidance. Thus, a mechanism for deeply integrating scene information into the motion generation process is necessary.
@@
Controllable Generation A ControlNet-like architecture is employed, where spatial control signals serve as inputs to the control branch for generating motions aligned with specified spatial trajectories, as illustrated in Figure 2. Key design elements consist of initializing the control branch with pretrained weights from the main branch and implementing zero-initialized connection layers. These design choices enable the incorporation of additional control capabilities while preserving the original model’s generative performance. Differ from (Xie et al., 2023), our proposed method achieves full-body control encompassing hand control. This advancement stems from optimized hand pose representation: wrist and 15 finger joints are parameterized using 6D global rotations within the world coordinate system, contrasting with SMPL-X’s rotations relative to parent joints. Deviations in trunk joint rotations result in cumulative error propagation through kinematic chains in conventional methods. The proposed representation determines hand poses through 16 global joint rotations combined with wrist positioning, effectively eliminating error accumulation pathways.
@@
Figure 4. Visulization of Guidance in SCoMoGen. Precise control of motions is achieved through classifier-free guidance. Green arrows represent joints being attracted by green anchors (hand joints, path waypoints). Red arrows represent the repulsive force on the joint (red body joint) close to the scene.
@@
Joint & Scene Guidance Despite the implemented spatial control mechanisms, generating motions that satisfy precision requirements remains challenging, since minor deviations in interactive tasks can produce noticeable motion artifacts. To address this limitation, classifier-free guidance (Ho and Salimans, 2022) is incorporated to facilitate human-scene interaction resolution without requiring additional motion capture annotations. As depicted in Figure 4, two complementary constraints joint-level and scene-level guidance are implemented within the framework.
@@
The joint constraint $\mathcal{L}_{j}$ establishes attraction forces between body joints $\mathcal{J}$ and target anchors $\mathcal{A}$ in 3D space. For root joint $j_{root}$ and path waypoints $a_{r}$, the loss is formulated as:
(6) $\mathcal{L}_{j}^{root}=\sum_{i=1}^{L_{r}}\|j_{r}^{(i)}-a_{r}^{(i)}\|_{2}^{2}$
@@
where $L_{r}$ is the number of path waypoints. For hand-object interactions, the constraint is implemented using::
(7) $\mathcal{L}_{j}^{hand}=\sum_{i=1}^{L_{h}}{\parallel j_{w}^{(i)}}-a_{w}^{\left(i\right)}\parallel_{2}^{2}+\sum_{i=1}^{L_{h}}{\parallel r_{h}^{(i)}}-a_{h}^{\left(i\right)}\parallel_{2}^{2}$
@@
where $j_{w}^{(i)}$ denotes wrist positions and $r_{h}^{(i)}$ represents 16 hand rotations. The parameter $L_{h}$ typically equals 2, corresponding to two key poses generated by SGAP for object pickup and placement. This value can be extended depending on application requirements.
@@
The scene constraint $\mathcal{L}_{s}$ generates adaptive repulsion forces through path-aligned point cloud contextual information. For each joint $j$, the repulsion loss is defined as:
(8) $\mathcal{L}_{s}=-\sum_{j\in\mathcal{J}}\sum_{p\in\mathcal{N}(j,\ 0.3\text{m})}\|j-p\|_{2}^{2}$
where $\mathcal{N}(j,r)$ represents scene points located within a radius $r$ of joint $j$. This formulation establishes proximity-based penalties while maintaining the original geometry characteristics.
Table 1. Quantitative results of human-object-scene interaction generation. This mainly involves the interaction between human and scenes, human and objects when characters operate objects in the scene. At the same time, this also evaluates the efficiency and accuracy of the character in carrying objects.

### v2

@@ Abstract
Generating high-fidelity full-body human interactions with dynamic objects and static scenes remains a critical challenge in computer graphics and animation. Existing methods for human-object interaction often neglect scene context, leading to implausible penetrations, while human-scene interaction approaches struggle to coordinate fine-grained manipulations with long-range navigation. To address these limitations, we propose HOSIG, a novel framework for synthesizing full-body interactions through hierarchical scene perception. Our method decouples the task into three key components: 1) a scene-aware grasp pose generator that ensures collision-free whole-body postures with precise hand-object contact by integrating local geometry constraints, 2) a heuristic navigation algorithm that autonomously plans obstacle-avoiding paths in complex indoor environments via compressed 2D floor maps and dual-component spatial reasoning, and 3) a scene-guided motion diffusion model that generates trajectory-controlled, full-body motions with finger-level accuracy by incorporating spatial anchors and dual-space gradient-based guidance. Extensive experiments on the TRUMANS dataset demonstrate superior performance over state-of-the-art methods. Notably, our framework supports unlimited motion length through autoregressive generation and requires minimal manual intervention. This work bridges the critical gap between scene-aware navigation and dexterous object manipulation, advancing the frontier of embodied interaction synthesis. Project page: https://yw0208.github.io/hosig/.
@@
Figure 1: Human-Object-Scene Interaction Generation. Our proposed HOSIG can generate high-fidelity full-body human motions. HOSIG can not only generate interactions with static scenes, but also generate object manipulation motions with fine hand-object contact. Moreover, relying on iterative generation and autonomous navigation, HOISG can generate long-term motions in complex indoor scenes.
Introduction
@@
Embodied intelligence represents a pivotal frontier in AI research, aiming to develop agents capable of navigating and manipulating 3D environments. While existing works have extensively explored human-object interaction (HOI) (Taheri et al. 2024; Zhang et al. 2024b; Petrov et al. 2024; Wu et al. 2022; Zheng et al. 2023; Li, Wu, and Liu 2023) and human-scene interaction (HSI) (Karunratanakul et al. 2023; Xiao et al. 2023; Zhao et al. 2023; Zhang and Tang 2022a; Wang et al. 2024b), few address the integrated human-object-scene interaction (HOSI) challenge (Wu et al. 2024; Jiang et al. 2024b, a; Lu et al. 2024). As shown in Figure 1, our goal is to enable characters to move in complex scenes, complete precise object operations, and seamlessly connect them into a long-term motion. The inherent complexity of synthesizing these multimodal interactions presents critical unsolved challenges in spatial reasoning and motion coordination.
@@
While recent advancements in HOI (Yang et al. 2024; Song et al. 2024; Peng et al. 2023; Taheri et al. 2022; Ghosh et al. 2023; Zhang et al. 2024a), they predominantly neglect 3D scene constraints. This oversight leads to implausible human-scene interpenetration, highlighting the necessity for scene-aware reasoning. Besides ignoring the scene in HOI, current HSI approaches exhibit two limitations: 1) Global scene encoding methods (Wang et al. 2024b, 2022b; Huang et al. 2023) lack granularity for precise motion synthesis, and 2) Local context perception strategies (Cen et al. 2024; Mao et al. 2022; Ghosh et al. 2021) struggle with pathfinding in complex environments, resulting in persistent interpenetration during long-range motion generation. These dual challenges underscore the need for unified scene-object-agent coordination.
@@
Our technical implementation features three key innovations. First, the grasp pose generator augments the cVAE framework (Taheri et al. 2022) with local scene geometry constraints, producing physically-plausible hand-object orientations that prevent scene interpenetration. Second, a novel 2D scene abstraction layer enables efficient navigation through 3D environments via heuristic pathfinding on compressed obstacle-aware maps, dynamically generating obstacle-aware motion graphs. Third, building upon ControlNet’s conditioning paradigm (Zhang, Rao, and Agrawala 2023), we develop a multi-condition diffusion framework that simultaneously integrates: (a) spatial anchors from navigation paths, (b) fine-grained hand control through grasp poses, and (c) path-aligned scene priors for continuous interaction optimization. This unified architecture achieves finger-level motion precision without auxiliary hand modules, surpassing previous trajectory-control methods through iterative spatial constraint.
@@
Related Work
Human-Object Interaction Generation
Research on human-object interaction (HOI) motion generation has progressed through two main paradigms. Early approaches primarily utilized reinforcement learning (RL), with initial works like (Peng et al. 2019, 2021) achieving basic interactions such as box touching. Subsequent RL methods developed more complex skills including basketball dribbling (Liu and Hodgins 2018), tennis playing (YUAN and Makoviychuk 2023), multi-object manipulation (Wang et al. 2023), and box carrying (Hassan et al. 2023).
The second paradigm employs generative models for direct motion synthesis. GOAL (Taheri et al. 2022) pioneered this direction using cVAEs for grasping motions. IMoS (Ghosh et al. 2023) later enabled action-label conditioned generation (e.g., photography, drinking). Recent works focus on language-conditioned generation: OOD-HOI (Zhang et al. 2024b) synthesizes human-object motions separately with contact optimization, while HOI-Diff (Peng et al. 2023) employs affordance-guided diffusion. Other key innovations include TriDi’s unified modeling of bodies/objects/contacts (Petrov et al. 2024) and CHOIS’s trajectory-constrained synthesis (Li et al. 2024b). InterDiff (Xu et al. 2023) further enables sequential generation from initial states.
Human-Scene Interaction Generation
@@
Figure 2: Overview of Our Pipeline. HOSIG can iteratively generate long-term motions based on spatial information, text, and the previous motion clip. There are three parts worth noting in the pipeline: (1) SGAP generates fine grasping postures to ensure the quality of character interaction. (2) Heuristic Navigation generates sparse human root joint trajectories to constrain the subsequent generated motions to be within the traversable area. (3) SCoMoGen uses a dual-branch design to achieve spatial control and adds additional joint & scene guidance during inference to achieve high-precision control.
For locomotion generation, existing methods adopt either data-driven (Wang et al. 2024a; Zhang and Tang 2022b; Rempe et al. 2023) or algorithmic approaches (Wang et al. 2022a). While data-driven methods struggle in novel complex environments, our work proposes an algorithmic solution for robust navigation. Semantic interaction synthesis initially produced static poses in scenes (Zhang et al. 2020; Li and Dai 2024; Zhao et al. 2022; Xuan et al. 2023), later evolving into language-guided systems (Wang et al. 2022b, 2024b; Huang et al. 2023) though requiring post-hoc optimization for scene compliance. Recent advances include concurrent navigation-interaction frameworks (Yi et al. 2024) and video-generation-based methods (Li et al. 2024a) with strong generalization but limited contact realism and motion range.
Closest to our approach, (Jiang et al. 2024b, a) enable triadic human-object-scene interactions. However, these require frame-level action labels and precise finger positions, whereas our method achieves comparable results using only initial/final object states.
Method
Problem Formulation
@@
Scene-Aware Grasp Pose Generation
@@
$X=\left[\varTheta,\beta,\mathbf{V},\mathbf{D}^{h\rightarrow o},\hat{\mathbf{h}},\mathbf{t}^{o},\mathbf{B}^{o},\mathbf{B}^{s}\right]$ (1)
@@
Figure 3: Visulization of Scene-Aware in SGAP. The green box is centered on the purple bottle. In SGAP, only the sparse scene point cloud inside the box is used, as shown in the right figure.
The critical component $\mathbf{B}^{s}$ constitutes the primary scene perception mechanism in SGAP. As illustrated in Figure 3, given a z-up object translation $\mathbf{t}^{o}=(x,y,z)$, we construct a scene context volume bounded by $x-0.8,x+0.8]\times[y-0.8,y+0.8]\times[0.2,1.8]$ meters, forming a $1.6^{3}\ m^{3}$ cubic region. To capture scene geometry, we employ a volumetric sampling strategy that generates a dense point cloud $\mathcal{P}_{s}$ containing both surface vertices and interior points from the scene mesh. This sampling follows a voxel grid resolution of 8 cm³, ensuring complete spatial coverage. The BPS transformation processes $\mathcal{P}_{s}$ through basis projections to produce the scene encoding $\mathbf{B}^{s}\in\mathbb{R}^{1024}$. These scene features condition the pose generation process, enabling synthesis of scene-adapted full-body poses.
@@
$\mathcal{L}_{sd}=-\frac{1}{N}\sum_{j=1}^{22}\sum_{k=1}^{N}\|\mathbf{J}_{j}-\mathbf{S}_{k}\|_{2}^{2}$ (2)
@@
where $\mathbf{J}_{j}\in\mathbb{R}^{3}$ denotes the $j$-th body joint position from the SMPL-X kinematic tree. $\mathbf{S}_{k}\in\mathbb{R}^{3}$ represents the $k$-th point in the localized scene point cloud $\mathcal{P}_{s}$, as shown in Figure 3. Point numbers $N$ is variable, depending on the number of point clouds in the current localized scene. This formulation imposes stronger penalties as joints approach scene surfaces. During training, $\mathcal{L}_{sd}$ backpropagates collision-avoidance constraints through the cVAE’s latent space, encouraging the generator to produce poses maintaining clearance from scene geometry while preserving natural motion kinematics.
@@
During inference, SGAP’s output parameters undergo selecting to extract three critical control signals: root joint translation $\mathbf{a}_{r}\in\mathbb{R}^{3}$, wrist joint position $\mathbf{a}_{w}\in\mathbb{R}^{3}$, and hand joint rotations $\mathbf{a}_{h}\in\mathbb{R}^{15\times 6}$. These elements constitute the spatial anchor tuple $\mathcal{A}=(\mathbf{a}_{r},\mathbf{a}_{w},\mathbf{a}_{h})$, which serves as key constraints for subsequent motion synthesis.
Heuristic Navigation on 2D Obstacle-Aware Map
Figure 4: Pipeline of Heuristic Navigation. The blue ball in the original scene is the starting point, and the red ball is the end point. The obstacle-aware map is presented in the form of a heat map, and the values correspond to the axis at the bottom. In Heuristic Function, the dark blue dot represents the current node, the light blue dot represents the candidate node, and the red star represents the end point.
Our Heuristic Navigation framework constitutes a customized A* variant specifically engineered for 3D point cloud environments, designed as a plug-and-play module compatible with diverse motion generation models. Crucially, the algorithm maintains full replaceability. Any trajectory synthesis method satisfying interface requirements can be integrated into our pipeline. This design intentionally validates our hypothesis: global scene perception substantially outperforms local perception strategies for long-range navigation in complex 3D environments. Here are the steps of our method:
As illustrated in Figure 4, the pipeline executes four cohesive stages. Initially, the 3D point cloud is partitioned into volumetric blocks and compressed vertically into a 2D heat map, i.e., obstacle-aware map, where each grid cell encodes the cumulative point numbers along the vertical axis. Then, this representation undergoes convolution smoothing to deliberately blur boundaries between obstacles and walkable regions, mitigating the algorithm’s tendency to generate trajectories adhering close to obstacle surfaces. Subsequently, a dense 2D path is computed via our implementation, employing a dual-component heuristic function. As shown in the lower left corner of the Figure 4, one term evaluates Euclidean distance to the goal, like the orange line. While the other term utilizes the Bresenham line algorithm to compute cumulative traversal costs by summing values along the direct path between candidate points and the destination, like the light orange area. Finally, the resulting 2D dense path is adaptively downsampled, added height, and transformed back into the native 3D coordinate system to yield a 3D sparse trajectory. This trajectory will be added to the $\mathbf{a}_{r}$ of the spatial anchor tuple $\mathcal{A}$ as a root joint control signal to serve the subsequent controllable motion generation.
Due to space limitations, if you are curious about the details of this algorithm, please refer to the detailed description in the supplementary materials.
Scene-Guided Controllable Motion Generation
@@
To enable the generation of high-fidelity interactions, two key challenges must be addressed. First, effective control over human body movement along the specified trajectory is required, with particular emphasis on enabling precise hand movements for object grasping. Second, in narrow and complex scenes, relying solely on the trajectory is insufficient to guarantee scene avoidance. Thus, a mechanism for deeply integrating scene information into the motion generation process is necessary. To address these two challenges, we respectively propose two corresponding approaches.
@@
Controllable Generation A ControlNet-like architecture is employed, where spatial control signals serve as inputs to the control branch for generating motions aligned with specified spatial trajectories, as illustrated in Figure 2. Specifically, the main branch of SCoMoGen takes the motion noises $\left\{\tilde{M}_{t}\right\}_{t=0}^{L}$ and the text embedding encoded by CLIP as inputs, and outputs next motion segmentation. Among them, for the coherence of motions, $\tilde{M}_{0}$ to $\tilde{M}_{10}$ are the last 11 frames of the previous motion segment. The control branch of SCoMoGen takes control signals as inputs. The control signals are converted from spatial anchor tuples $\mathcal{A}$, where the uncontrolled parts are set to 0. Key design elements consist of initializing the control branch with pretrained weights from the main branch and implementing zero-initialized connection layers. These design enable the incorporation of additional control capabilities while preserving the original model’s generative performance. Notably, for the first time, our proposed SCoMoGen achieves full-body control encompassing hand control. This advancement stems from optimized hand pose representation: wrist and 15 finger joints are parameterized using 6D global rotations within the world coordinate system, contrasting with SMPL-X’s rotations relative to parent joints. Deviations in trunk joint rotations result in cumulative error propagation through kinematic chains in conventional methods. The proposed representation determines hand poses through 16 global joint rotations combined with wrist positioning, effectively eliminating error accumulation pathways.
@@
Figure 5: Visulization of Guidance in SCoMoGen. Precise control of motions is achieved through gradient-based guidance. Green arrows represent joints being attracted by green anchors (hand joints, path waypoints). Red arrows represent the repulsive force on the joint (red body joint) close to the scene.
@@
Joint & Scene Guidance Despite the implemented spatial control mechanisms, generating motions that satisfy precision requirements remains challenging, since minor deviations in interactive tasks can produce noticeable motion artifacts. To address this limitation, gradient-based guidance (Guo et al. 2024) is incorporated to facilitate human-scene interaction generation without requiring additional motion post-processing. As depicted in Figure 5, two complementary constraints joint-level and scene-level guidance are implemented within the framework.
@@
The joint constraint $\mathcal{L}_{j}$ establishes attraction forces between body joints $\mathcal{J}$ and target anchors $\mathcal{A}=(\mathbf{a}_{r},\mathbf{a}_{w},\mathbf{a}_{h})$ in 3D space. For root joint $j_{r}$ and root anchors $\mathbf{a}_{r}$, the loss is formulated as:
$\mathcal{L}_{j}^{r}=\sum_{i=1}^{L_{r}}\|j_{r}^{(i)}-\mathbf{a}_{r}^{(i)}\|_{2}^{2}$ (3)
@@
where $L_{r}$ is the number of path waypoints. For hand-object interactions, the constraint is implemented using:
$\mathcal{L}_{j}^{hand}=\sum_{i=1}^{L_{h}}{\parallel j_{w}^{(i)}}-\mathbf{a}_{w}^{\left(i\right)}\parallel_{2}^{2}+\sum_{i=1}^{L_{h}}{\parallel r_{h}^{(i)}}-\mathbf{a}_{h}^{\left(i\right)}\parallel_{2}^{2}$ (4)
@@
where $j_{w}^{(i)}$ denotes wrist positions and $r_{h}^{(i)}$ represents 16 hand rotations. The parameter $L_{h}$ typically equals 2, corresponding to two key poses generated by SGAP for object pickup and placement. This parameter can be extended depending on application requirements.
@@
The scene constraint $\mathcal{L}_{s}$ generates adaptive repulsion forces through path-aligned point cloud. For each joint $j$, the repulsion loss is defined as:
$\mathcal{L}_{s}=-\sum_{j\in\mathcal{J}}\sum_{p\in\mathcal{N}(j,\ 0.3\text{m})}\|j-p\|_{2}^{2}$ (5)
where $\mathcal{N}(j,r)$ represents scene points located within a radius $r$ of joint $j$. By restricting attention to the local $\mathcal{N}(j,r)$, irrelevant scene elements are excluded, sharply reducing computational overhead and steering optimization toward a more precise direction.